OpenStudy (anonymous):

sue can shovel snow from her driveway in 50 mins. tom can do the same job in 40 mins. how long would it take sue and tom to do the job if they worked together

OpenStudy (anonymous):

why do you have the name dumbcow out of curiosity

OpenStudy (dumbcow):

inside joke

OpenStudy (anonymous):

i see

OpenStudy (dumbcow):

can never think of a cool name, so i just went with this one

OpenStudy (anonymous):

mine has been my screenname for over 20 yrs lol

OpenStudy (dumbcow):

1/50 + 1/40 = 4/200 + 5/200 = 9/200 9/200 t = 1 t = 200/9 = 22.22 min

OpenStudy (anonymous):

1/{1/40+1/50}

OpenStudy (anonymous):

200/90=22.22min

OpenStudy (anonymous):

panthor got it??

OpenStudy (anonymous):

yeah ty